andrew44234:

A graph is shown below: dotted line joining ordered pair negative 1, negative 2 and 5, 0 and the region above this line is shaded Which of the following inequalities is best represented by this graph?

charge69:

I'd suggest that you graph the scenario to get a better understanding or at least make a rough sketch if you are unable to graph. It's dotted, so it's either between < or >. {(-1,-2),(5,0)} is the set of points, yielding m=1/3 and a y-intercept of b=-5/3. Since it's shaded above the line, the inequality is > So the inequality is y>(1/3)x-5/3